OpenStudy (anonymous):

Can someone explain why \[\int\limits x ^{-1} dx\] is anti-derivative \[\int\limits \ln(-x) + C\]

OpenStudy (anonymous):

it isn't

OpenStudy (anonymous):

meaning?

OpenStudy (anonymous):

\[\int \frac{dx}{x}=\ln(x)\]

OpenStudy (anonymous):

well \[x^{-1} = 1/x\]?

OpenStudy (anonymous):

\[\ln(-x)=\ln(\frac{1}{x})\] although i don't know if that helps any the anti - derivative of \(x^{-1}\) is \(\ln(x)\)

OpenStudy (anonymous):

i am confused by your question, because you have two indefinite integrals in it

OpenStudy (anonymous):

how about this can you work out first equation

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

plz <3

OpenStudy (anonymous):

ok what is it?

OpenStudy (anonymous):

i don't get why if you use \[\frac{ x ^{n+1}}{ n+1 }\] that it would not just be \[x^0 + C\]

OpenStudy (anonymous):

ok i think i understand the question it is this: what is\[\int x^{-1}dx\] or equivalently what is \[\int \frac{1}{x}dx\] right?

OpenStudy (anonymous):

since the exponent is \(-1\) why can't you just use the power rule backwards, i.e. us \[\int x^ndx=\frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

it doesn't work in the case \(x^{-1}\) for a couple reasons, but one is this: if you used it literally you would get \[\int x^{-1}dx=\frac{x^{-1+1}}{-1+1}=\frac{x^0}{0}\]

OpenStudy (anonymous):

and of course you cannot divide by zero!

OpenStudy (anonymous):

so that's is why i cant use the power rule backwards right?

OpenStudy (anonymous):

yes, it is not applicable in the this case because the denominator is zero

OpenStudy (anonymous):

now as to why \[\int x^{-1}dx=\ln(x)\], that depends largely on your definition of the log but one simple reason is that \[\frac{d}{dx}[\ln(x)]=\frac{1}{x}\] so pretty much by definition, \[\int \frac{1}{x}dx=\ln(x)\]

OpenStudy (anonymous):

later on in this or some other math class you will DEFINE the log to be \[\ln(x)=\int_1^x\frac{dt}{t}\]

OpenStudy (anonymous):

at the moment my guess is your definition of the log is the "inverse of the exponential"

OpenStudy (anonymous):

i sorta get it.

OpenStudy (anonymous):

the derivative of \(\ln(x)\) is \(\frac{1}{x}\) so the anti derivative of \(\frac{1}{x}\) is \(\ln(x)\) that should work for your calc class

OpenStudy (anonymous):

i am assuming you know that the derivative of \(\ln(x)\) is \(\frac{1}{x}\) from before, rigth?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i was just wondering why it would not be x^0 but you explain that very well

OpenStudy (anonymous):

ok then that explanation should be sufficient, because if the derivative of \(F\) is \(f\) then the anti derivative of \(f\) is \(F\)

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

alright man thanks for the help

OpenStudy (anonymous):

yw